OpenStudy (anonymous):

Let a,b,c be positive real numbers. If the minimum value of \(\Large \frac{(b+c-a)^2}{(b+c)^2+a^2} + \frac{(c+a-b)^2}{(c+a)^2+b^2} + \frac{(a+b-c)^2}{(a+b)^2+c^2} \) can be written as \(\Large \frac{m}{n}\), then find \(\Large m+n\)

OpenStudy (anonymous):

I think we should use cauchy-schwarz inequality

ganeshie8 (ganeshie8):

Clearly this is symmetric in a,b,c so the min/max occurs only when a=b=c

ganeshie8 (ganeshie8):

symmetry argument may not look pleasing for everybody, lets see if we can work it using cauchy-schwarz inequality

OpenStudy (anonymous):

using cauchy-schwarz inequality on each of these pieces gives \(\Large \frac{(b+c-a)^2}{(b+c)^2+a^2} + \frac{(c+a-b)^2}{(c+a)^2+b^2} + \frac{(a+b-c)^2}{(a+b)^2+c^2}\\\Large \le2 [\frac{(b+c)^2-a^2}{(b+c)^2+a^2}+\frac{(c+a)^2-b^2}{(c+a)^2+b^2}+\frac{(a+b)^2-c^2}{(a+b)^2+c^2}] \)

OpenStudy (anonymous):

so, what is the next step?

OpenStudy (anonymous):

@vishweshshrimali5

OpenStudy (vishweshshrimali5):

One minute I would have to go through all of this ^^^

OpenStudy (anonymous):

ok :)

OpenStudy (vishweshshrimali5):

Well @ganeshie8 's method is of course correct. This is a very simple method and works almost all times

OpenStudy (vishweshshrimali5):

Are you sure you applied the inequality correctly ?

OpenStudy (vishweshshrimali5):

Just conforming ...

OpenStudy (anonymous):

we guess is the minimum?

OpenStudy (vishweshshrimali5):

Appears so

OpenStudy (vishweshshrimali5):

I am on my way home and don't have any pen/copy. So, I would let you know the solution when I reach home :)

OpenStudy (anonymous):

okay, I wait

OpenStudy (vishweshshrimali5):

Sure :)

OpenStudy (vishweshshrimali5):

Okay I have not reached home yet but I have an idea

OpenStudy (vishweshshrimali5):

We assume that 3/5 is the minimum value

OpenStudy (vishweshshrimali5):

I can write the given expression in summation form as: \[\large{\sum {\cfrac{(b+c-a)^2}{(b+c)^2+a^2}}}\] And I have assumed that: \[\large{\sum {\cfrac{(b+c-a)^2}{(b+c)^2+a^2}} \ge \cfrac{3}{5}}\]

OpenStudy (vishweshshrimali5):

*I will have to prove this ^^^

OpenStudy (vishweshshrimali5):

Now I am going to subtract 1 from each term, and thus subtracting 3 from 3/5

OpenStudy (vishweshshrimali5):

I will get: \[\large{\sum ({\cfrac{(b+c-a)^2}{(b+c)^2+a^2}}) - 1 \ge \cfrac{3}{5} -3}\] \[\large{\implies \sum (\cfrac{(b+c-a)^2 - (b+c)^2 - a^2}{(b+c)^2+a^2}) \ge \cfrac{-12}{5}}\] \[\large{\implies \sum (\cfrac{-2ab - 2ac}{(b+c)^2+a^2}) \ge \cfrac{-12}{5}}\] \[\large{\implies \sum \cfrac{2ab+2ac}{(b+c)^2+a^2} \le \cfrac{12}{5}}\]

OpenStudy (vishweshshrimali5):

Okay I can think of only one thing: Add 3 on both sides. \[\large{\implies \sum (\cfrac{2ab+2ac}{(b+c)^2+a^2} +1) \le \cfrac{12}{5} + 3}\] \[\large{\implies \sum (\cfrac{2a(b +c) + (b+c)^2 + a^2}{(b+c)^2+a^2}) \le \cfrac{27}{5}}\] \[\large{\implies \sum \cfrac{(a+b+c)^2}{a^2+(b+c)^2} \le \cfrac{27}{5}}\] \[\large{\implies (a+b+c)^2 \sum \cfrac{1}{a^2+(b+c)^2} \le \cfrac{27}{5}}\]

OpenStudy (vishweshshrimali5):

Okay, I would have to think something more for further steps :)

OpenStudy (vishweshshrimali5):

@ganeshie8 any idea ?

OpenStudy (vishweshshrimali5):

@JungHyunRan what do you think ?

OpenStudy (vishweshshrimali5):

Okay since the inequality is homogeneous, I can normalize the inequality. Yup this may work :)

OpenStudy (vishweshshrimali5):

Without loss of generality, I can assume that: \[\large{a+b+c = 3}\] This will help me in simplifying the fraction in RHS into 3/5.

OpenStudy (vishweshshrimali5):

So, I get, \[\large{\sum \cfrac{1}{(3-a)^2 + a^2} \le \cfrac{3}{5}}\]

OpenStudy (vishweshshrimali5):

Any doubt till this step @JungHyunRan ?

OpenStudy (vishweshshrimali5):

I tried using CS inequality but it didn't prove much beneficial

OpenStudy (anonymous):

do we have general solution for this type ?

OpenStudy (vishweshshrimali5):

Well for different inequalities there are different methods.

OpenStudy (vishweshshrimali5):

I tried to use Jensen's inequality at this step but since the function is convex I cannot get the maximum value using this inequality. So, lets move on

OpenStudy (vishweshshrimali5):

Okay I can write the last inequality as: \[\large{\sum \cfrac{1}{(3-a)^2 + a^2} \le \cfrac{3}{5}}\] \[\large{\implies \sum \cfrac{5}{(3-a)^2+a^2} \le 3}\] \[\large{\implies \sum (\cfrac{5}{(3-a)^2 + a^2} -1) \le 0}\] \[\large{\implies \sum (1-\cfrac{5}{(3-a)^2+a^2}) \ge 0}\] Okay stuck again :P

OpenStudy (vishweshshrimali5):

I have to prove my assumption.

OpenStudy (vishweshshrimali5):

So, I am basically moving backwards. I assumed that the minimum value is 3/5. Now, to prove my assumption, I try to use this minimum value to get a step which is true because of given information in problem. If I can arrive at one such step, then I would be able to prove that my assumption was correct and that 3/5 is the minimum value.

OpenStudy (anonymous):

i see

OpenStudy (vishweshshrimali5):

Okay lets see if this works out: \[\large{\sum 1 - \cfrac{5}{(3-a)^2+a^2}}\] \[\large{= \sum 1 - \cfrac{5}{9+2a^2-6a}}\] Now, I got 3/5 in the very first step by assuming that: \[\large{a=b=c}\] Then, in a later step during normalization I assumed that: \[\large{a+b+c = 3}\] Thus, I can write that I have assume: \[\large{a = b= c =1}\]

OpenStudy (vishweshshrimali5):

No that's not correct

OpenStudy (vishweshshrimali5):

\(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 Okay lets see if this works out: \[\large{\sum 1 - \cfrac{5}{(3-a)^2+a^2}}\] \[\large{= \sum 1 - \cfrac{5}{9+2a^2-6a}}\] Now, I got 3/5 in the very first step by assuming that: \[\large{a=b=c}\] Then, in a later step during normalization I assumed that: \[\large{a+b+c = 3}\] Thus, I can write that I have assume: \[\large{a = b= c =1}\] \[\color{red}{\text{FORGET IT!!}}\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (vishweshshrimali5):

This is what that happens whenever I try to apply more knowledge than necessary :P

OpenStudy (anonymous):

^^

OpenStudy (vishweshshrimali5):

Yeah... Okay may be something else may help ... Will have to think :)

OpenStudy (anonymous):

you are so awesome, so smart :D

OpenStudy (vishweshshrimali5):

Okay: \[\large{\sum(1-\cfrac{5}{2a^2-6a+9})}\] \[\large{=2(a^2-3a+2)\sum\cfrac{1}{2a^2-6a+9}}\] \[\large{=2(1-a)(2-a)\sum\cfrac{1}{2a^2-6a+9}}\] \[\large{=2(1-a)\sum \cfrac{-5a+10}{5(2a^2-6a+9)}}\] \[\large{=2(1-a)\sum \cfrac{-5a+10}{5(2a^2-6a+9)}}\] \[\large{=2(1-a)\sum \cfrac{-5a+10}{5(2a^2-6a+9)}}\] \[\large{=2(1-a)\sum \cfrac{2a^2-6a+9 + a + 1 -2a^2}{5(2a^2-6a+9)}}\] \[\large{=2(1-a)\sum \cfrac{2a^2-6a+9 + (a+1-2a^2)}{5(2a^2-6a+9)}}\] \[\large{=2(1-a)\sum \cfrac{2a^2-6a+9 + (1-a)(2a+1)}{5(2a^2-6a+9)}}\] \[\large{=\sum \cfrac{2(1-a)(2a^2-6a+9) + 2(a-1)^2(2a+1)}{5(2a^2-6a+9)}}\] \[\large{=\sum \cfrac{2(1-a)}{5} + \cfrac{2(a-1)^2(2a+1)}{5(2a^2-6a+9)}}\]

OpenStudy (vishweshshrimali5):

\[\large{\ge \sum \cfrac{2(1-a)}{5}}\] \[\large{=0}\] Thus, \[\large{\sum(1-\cfrac{5}{2a^2-6a+9}) \ge 0}\] is true. Hence, \[\large{\sum \cfrac{1}{(3-a)^2 + a^2} \le \cfrac{3}{5}}\] is true Hence, 3/5 is the minimum value of the given expression. Thus, m+n = 8

OpenStudy (vishweshshrimali5):

And thanks @JungHyunRan

OpenStudy (vishweshshrimali5):

@ganeshie8 @JungHyunRan I would be offline for 30 - 45 minutes till then can you please check out the solution

OpenStudy (anonymous):

ok :)

OpenStudy (vishweshshrimali5):

I am back! I am back ! :D

OpenStudy (vishweshshrimali5):

Any doubt @JungHyunRan in the solution ?

OpenStudy (anonymous):

no, :) I see the solution